# Rheologically tuned modes of collective transport in active viscoelastic films

### v1

@@ Abstract
Many living biological materials simultaneously exhibit properties of both viscous fluids and elastic solids [1] or may even change their behavior dynamically. For example, during biofilm formation, bacteria excrete extracellular polymeric substances [2, 3], which results in complex rheological behavior [4]. Depending on the stage of biofilm formation, the location within it, and the time scales considered, the material is rather fluid- or solid-like or somewhere in between. From a physical perspective, such biological systems are described within the framework of active matter [5, 6]. It concerns the collective behavior of active, self-propelled agents, frequently embedded in a surrounding medium. However, most studies refer to one of two categories, depending on the considered long-term flow behavior: active fluids or active solids.
@@
Active fluids lack long-term elastic restoring forces and are capable of terminal flow. Primary examples are fluid suspensions of microswimmers [7, 8], such as bacteria or artificial self-propelled particles, and active nematics [9]. As described by the seminal Vicsek model [10] and its continuum counterpart, the Toner–Tu theory [11, 12], collections of active agents capable of unrestricted long-term movement may develop polar orientational order and directed collective motion. In addition to these flocking or swarming states [13, 14], active fluids exhibit various spatiotemporal patterns [15, 16, 17] including active turbulence [18, 19, 20, 21].
@@
In contrast to active fluids, active solids are unable to support terminal flow due to elastic restoring forces. Instead, they exhibit reversible elastic deformations. In theoretical descriptions, the continuous nature of the embedding medium is often reduced to elastic spring networks of connected active agents [22, 23, 24, 25, 26]. Other examples of active solids are driven by contractile or extensile active elements and resulting stresses within the medium [27, 28], as in muscular tissue [29]. Recent topics of interest in active solids include the nature of excitations [30, 31] and the phenomenon of odd elasticity [32, 33].
@@
There are numerous theoretical studies on either active solid-like [34, 28, 32] or active fluid-like materials with simple [12, 35, 20] or complex [36, 37, 38, 39, 40] rheologies. However, a comprehensive exploration of the spatiotemporal dynamics covering these two limiting cases and the broad spectrum in between is yet to be presented.
@@
We aim to bridge this gap and focus on the interplay of viscous and elastic properties in active media on substrates. To this end, we employ a unified description that can be tuned continuously from the limit of active fluid-like to active solid-like properties. This framework, which covers a broad range of viscoelastic behavior in between, is derived in a recent article [41]. It is motivated from a generic microscopic picture of active agents and combined with a memory-based continuum description of the embedding medium. Basic, spatially uniform analytical solutions for resulting dynamic states were obtained for small idealized systems.
@@
Here, we leave this idealization and present a comprehensive dynamical state diagram. It reveals fascinating and novel dynamic states when varying the activity of the suspension and the rheological properties from fluid-, via viscoelastic, to solid-like. Numerical calculations uncover complex spatiotemporal states, including stripe-like patterns in the active viscoelastic regime and separated oscillating domains for increasingly solid-like systems. Moreover, we illustrate how tuning the rheological properties in such active suspensions can be exploited for active collective transport on the microscopic scale. It allows for intermediate storage times, maintaining the spatial arrangement within a set of cargo units.
@@
In short, in the theoretical framework, the behavior is governed by the interplay of three continuum fields [41]. The coarse-grained polar order parameter field $\mathbf{P}$ characterizes the local alignment of active units. It couples to the overall velocity field $\mathbf{v}$ and to the elastic displacement field $\mathbf{u}$. The latter captures the deformation history of the film.
@@
where $\mathbf{I}$ is the unit matrix. In addition to advection via the flow field $\mathbf{v}$, Eq. (1) includes reorientation via the vorticity tensor $\boldsymbol{\Omega}=[(\nabla\mathbf{v})^{\top}-(\nabla\mathbf{v})]$, where denotes the transpose. The first two terms on the right-hand side arise due to orientational and translational diffusion, while the diffusion coefficients were used to rescaled time and space. The term proportional to the alignment strength $\gamma_{\mathrm{a}}$ captures the tendency of active units to align with the overall flow field $\mathbf{v}$ near a supporting substrate [42, 43, 44, 45]. Similar terms are included in various discretized models of active solids [46, 47, 30, 48]. Nonlinearity limits the magnitude of the polar order parameter to $|\mathbf{P}|\leq 1$.
@@
We here assume an incompressible material, $\nabla\cdot\mathbf{v}=0$, enforced by the pressure field $p$. Viscosity $\eta$ and shear modulus $\mu$ together with the displacement field $\mathbf{u}$ quantify viscous and elastic forces. Underlying friction with the substrate is set by $\nu_{\mathrm{v}}$ and slows down movement. Similarly, elastic restoring forces quantified by $\nu_{\mathrm{d}}$ result from interactions with the substrate and pull material elements back to their memorized positions. Active forcing drives the dynamics of the film and is set by $\nu_{\mathrm{p}}$.
Third, the displacement field is governed by
@@
While on the right-hand side overall motion $\mathbf{v}$ generates distortions, the memory of displacement $\mathbf{u}$ decays with a relaxation time $\tau_{\mathrm{d}}$. This relaxation process implies a net flow and thus incomplete elasticity. Full elastic reversibility for solids is found in the limit $\tau_{\mathrm{d}}\to\infty$. Finite values $\tau_{\mathrm{d}}$ describe a broad range of viscoelastic fluids, while the limit $\tau_{\mathrm{d}}\to 0$ refers to purely viscous fluids. Thus, rheology can be tuned continuously from fluid- to solid-like behavior. Equation (3) includes convective transport $\mathbf{v}\cdot\nabla\mathbf{u}$ by overall flow. In combination with Eq. (2), it represents a first step beyond linear elasticity [49]. This description can be derived from a generic continuum theory [50] based purely on conservation laws and symmetry arguments upon linearization in $\nabla\mathbf{u}$ [41]. The contribution $-\nabla q$ there arises from incompressibility $\nabla\cdot\mathbf{u}=0$ under linearized strains [41].
Figure 1: Dynamical state diagram as a function of the displacement relaxation time $\tau_{\mathrm{d}}$ and strength of active forcing $\nu_{\mathrm{p}}$. From left to right, the influence of elasticity increases. Colors indicate regions of different stable dynamical states. Insets in the corners illustrate idealized tracer trajectories. Blue and red dots refer to the dynamical states explored in more detail in Figs. 2 and 3, respectively. Remaining parameter values are set to $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, and $\nu_{\mathrm{d}}=10$.
Previous derivation of the theory suggested three analytical, idealized, spatially uniform mathematical solutions [41]. We here challenge their stability, amended by numerical simulations. For technical details see the Supplemental Material [51], which includes Refs. 52, 53, 54. The central goal is to construct the dynamical state diagram for active films of viscoelastic rheology.
@@
Raising the influence of elasticity by increasing $\tau_{\mathrm{d}}$, we find that both states extend to the right in Fig. 1. Yet, elasticity seems to hinder motion. Illustratively, the elastic restoring forces pin the active agents. Therefore, the phase boundary between polar moving (green) and isotropic resting (yellow) areas turns towards higher activity. The phase boundary results from linear stability analysis and is confirmed numerically [51].
The picture changes notably when we increase elasticity at elevated levels of activity from the polar state. Then novel, qualitatively different behavior emerges. It is marked by the blue region in Fig. 1. Numerical evaluations indicate a stripe-like dynamic state, see Fig. 2(a). The inset depicts the associated wiggling, snake-like motion.
@@
Indeed, the state of globally uniform polar orientational order becomes linearly unstable in this region [51]. The growth rate $\mathrm{Re}\{\lambda(k=|\mathbf{k}|)\}$ is plotted in Fig. 2(b), with wavevectors $\mathbf{k}$ oriented parallel to $\mathbf{P}$. Importantly, the fastest-growing mode is of finite wavenumber $k_{\mathrm{max}}>0$. Thus, the instability features a certain length scale $\Lambda=2\pi/k_{\mathrm{max}}$, at least close to its onset. This length is reflected by spatial modulations along the migration direction, see Fig. 2(a). There, the local velocity field is indicated by the dark arrows. They deviate periodically in space to the left and to the right of the average migration direction given by $\langle\mathbf{v}\rangle$. The periodic modulation is further visualized when we calculate the vorticity field $\omega=(\nabla\times\mathbf{v})_{z}$, see the colored stripe pattern in Fig. 2(a).
@@
Moreover, from $\mathrm{Im}\{\lambda(k)\}$, we calculate the phase velocity of the pattern, $v_{\mathrm{ph}}=-\mathrm{Im}(\lambda)/k$. It characterizes the traveling speed of the individual waves. Interestingly, we always find $v_{\mathrm{ph}}<0$. Consequently, the emerging structures travel in the direction opposite to global motion $\langle\mathbf{v}\rangle$, see Fig. 2(a). Supplemental Movie 1 provides a dynamic illustration.
Further increasing the influence of elasticity via $\tau_{\mathrm{d}}$ to the right in Fig. 1, we cross the red line. It marks the existence of a basic analytical solution, corresponding to rotating spatially uniform polar order $\mathbf{P}$, flow $\mathbf{v}$, and displacement $\mathbf{u}$ [51]. Yet, we do not observe this state in our spatially extended numerical systems.
Here, numerically, we identify a hysteretic regime until we reach the dotted line in Fig. 1 [51]. In this area, the emergent state depends sensitively on the initial conditions. On the one hand, upon initialization by a weakly perturbed state of global polar orientational order, still the stripe patterns develop. On the other hand, upon initialization by a weak perturbation of the above-mentioned analytical solution [51], rotations of the three fields are maintained. Further to the right of the dotted line in Fig. 1, the stripes are not found as a final dynamic state any more. Although transients can be observed for different initializations, at the end, we find a state of locally rotating fields $\mathbf{P}$, $\mathbf{v}$, and $\mathbf{u}$, see Fig. 3(a). This state is not uniform, as the system splits up into different domains of rotation. Thus, it becomes spatially heterogeneous.
Physically, the elastic restoring forces for elevated $\tau_{\mathrm{d}}$ become too strong to allow for persistent motion. The globally ordered polar or stripe-like dynamic states are associated with continuous directed transport. However, this is suppressed by higher levels of elasticity. Instead, local rotations of the polar order $\mathbf{P}$, flow field $\mathbf{v}$, and displacement field $\mathbf{u}$ combine active transport with local elastic anchoring. Imperfect circular trajectories arise, see the top right insets in Figs. 1 and 3(a).
@@
Figure 3: Heterogeneous state of imperfect local circling motion. (a) At each spot, the vectors of polar orientational order $\mathbf{P}$ and collective flow $\mathbf{v}$ rotate over time. Dark arrows indicate $\mathbf{v}$. Domains are identified by the field of phase shift $\Delta\phi_{Pv}$ between $\mathbf{P}$ and $\mathbf{v}$, see the color bar. Positive and negative values are connected to clockwise and counterclockwise rotation, respectively. A sample trajectory is shown in the top right corner. The size of the snapshot is $100\times 100$. (b) Averaged spatial correlation function $C_{\Delta\phi_{Pv}}(\Delta x)$ of the phase shift $\Delta\phi_{Pv}$ between any two points of evaluation with distance $\Delta x$ from each other. The associated characteristic length scale $\ell$ of the dynamic patterns is identified with the first minimum of $C_{\Delta\phi_{Pv}}(\Delta x)$. Inset: Typical domain size $\ell$ as a function of active force strength $\nu_{\mathrm{p}}$ for different values of the relaxation time $\tau_{\mathrm{d}}$. Remaining parameter values are $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, $\nu_{\mathrm{d}}=10$, and $\nu_{\mathrm{p}}=20$.
@@
Here, $\Delta x$ is the distance between any two points in space and $\hat{\mathbf{n}}$ is the direction of their separation. $\langle\dots\rangle_{\hat{\mathbf{n}},\mathbf{x},t}$ denotes the average over orientations $\hat{\mathbf{n}}$, space positions $\mathbf{x}$, and time $t$. For the dynamically evolving state depicted in Fig. 3(a), we plot $C_{\Delta\phi_{Pv}}(\Delta x)$ in Fig. 3(b). The distinct length scale $\ell$ quantifying the typical domain size is identified with the location of the first minimum of $C_{\Delta\phi_{pv}}(\Delta x)$. Indeed, we find that the length scale $\ell$ of the typical size of the correlated domains is of the same order as the thickness of the stripes $\Lambda/2$ referred to in Fig. 2(a). The inset of Fig. 3(b) shows $\ell$ as a function of the strength of active forcing $\nu_{\mathrm{p}}$ for different values of the relaxation time $\tau_{\mathrm{d}}$ within the red region of Fig. 1. Mainly, we find that $\ell$ obtains a minimum, but then increases continuously with increasing activity $\nu_{\mathrm{p}}$. Thus, the structures become increasingly correlated. Approaching the transition to the isotropic state, see below, $\ell$ diverges and the structures are highly correlated in space. Overall, we find that elasticity in terms of the relaxation time $\tau_{\mathrm{d}}$ has only little influence on $\ell$ for the considered parameters.
@@
As an immediate perspective, changing the rheology of active films allows to control collective transport. For example, light and external electric fields enable remote dynamic tuning of the rheology of photorheological [55, 56, 57] or electrorheological [58] materials. Exploiting such effects facilitates externally controlled transport of cargo objects by active carrier media. Persistent polar orientational order implies directed transport, while switching to the nonpolar states facilitates intermediate storage times. Elevated activity implies increased speed of transport. It requires intermediate switching to the state of imperfect local circling during rest intervals. We numerically demonstrate this scenario in Fig. 4 by periodically changing the relaxation time $\tau_{\mathrm{d}}$, see Fig. 4(a). Resulting trajectories in Fig. 4(b) show the transport of an array of cargo objects, including intermediate storage periods, while maintaining the internal positional order within the array. Suitable preparation of the substrate can guide the transport. The approach opens new routes in microfluidics powered by the activity of microswimmers such as bacteria [59, 60, 61, 62].
In summary, we uncover a rich world of dynamic states that result from the interplay between viscoelasticity and activity. We construct the overall dynamical state diagram. For example, novel phases of active matter emerge in the form of stripe patterns and heterogeneous states of imperfect circling as an immediate consequence of activity and viscoelasticity. Several features are neither observed in purely viscous active fluids nor in purely elastic active solids.
Besides the quantitative characterization of engineered active systems to control directed transport of cargo on the microscale, our description relates to the complex dynamical behavior in living biological matter. Prime examples are bacterial biofilms [2, 3, 4]. They may change their rheological characteristics dynamically and continuously from purely viscous via viscoelastic to elastic as a function of time and position within the film.
@@
Figure 4: Rheologically controlled transport of an array of cargo objects. (a) We periodically vary the relaxation time $\tau_{\mathrm{d}}$ of elastic displacements between $0.25$ and $1$. Thus, we tune the degree of elasticity. (b) Resulting trajectories for an array of four transported cargo objects. Tuning rheology allows in a controlled way to switch between persistent collective transport and intermediate storage times, roughly maintaining the spatial order within the transported cargo array. Blue and red parts in (a) and (b) correspond to each other and to the blue and red regions in the state diagram in Fig. 1. The remaining parameters are set to $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, $\nu_{\mathrm{d}}=10$, and $\nu_{\mathrm{p}}=20$. The initial edges in the cargo array are of length $1$.

### v2

@@ Abstract
Many living biological materials simultaneously exhibit properties of both viscous fluids and elastic solids [1] or may even change their behavior dynamically. For example, during biofilm formation, bacteria excrete extracellular polymeric substances [2, 3], which results in complex rheological behavior [4]. Depending on the stage of biofilm formation, the location inside, and the considered time scales, the material is rather fluid- or solid-like or somewhere in between. From a physical perspective, such biological systems are described as active matter [5, 6]. It concerns the collective behavior of active, self-propelled agents, frequently embedded in a surrounding medium. However, most studies refer to one of two categories, depending on the considered long-term flow behavior: active fluids or active solids.
@@
Active fluids lack long-term elastic restoring forces and are capable of terminal flow. Primary examples are fluid suspensions of microswimmers [7, 8], such as bacteria or artificial self-propelled particles, and active nematics [9]. As described by the seminal Vicsek model [10] and its continuum counterpart, the Toner–Tu theory [11, 12], collections of active agents capable of unrestricted long-term movement may develop polar orientational order and directed collective motion. In addition to these flocking or swarming states [13, 14], active fluids exhibit various spatiotemporal patterns [15, 16, 17, 18] including active turbulence [19, 20, 21, 22].
@@
In contrast to active fluids, active solids are unable to support terminal flow due to elastic restoring forces. Instead, they exhibit reversible elastic deformations. In theoretical descriptions, the continuous nature of the embedding medium is often reduced to elastic spring networks of connected active agents [23, 24, 25, 26, 27]. Other examples are driven by contractile or extensile active elements and resulting stresses within the medium [28, 29], as in muscular tissue [30]. Recent topics of interest in active solids include the nature of excitations [31, 32] and the phenomenon of odd elasticity [33, 34].
@@
There are numerous theoretical studies on either active solid-like [35, 29, 33] or active fluid-like materials with simple [12, 36, 21] or complex [37, 38, 39, 40, 41, 42, 43] rheologies. However, a comprehensive exploration of the spatiotemporal dynamics covering these two limiting cases and the broad spectrum in between is yet to be presented.
@@
We aim to bridge this gap and focus on the interplay of viscous and elastic properties in active media on substrates. To this end, we employ a unified description that can be tuned continuously from the limit of active fluid-like to active solid-like properties. This framework, which covers a broad range of viscoelastic behavior in between, is derived in a companion article [44]. It is motivated from a generic microscopic picture of active agents and combined with a memory-based continuum description of the embedding medium. Basic, spatially uniform analytical solutions for resulting dynamic states were obtained for small idealized systems.
@@
Here, we leave this idealization and present a comprehensive dynamical state diagram. It reveals fascinating and novel dynamic states when varying the activity of the suspension and the rheological properties from fluid-, via viscoelastic, to solid-like. Numerical calculations uncover complex spatiotemporal states, including stripe-like patterns in the active viscoelastic regime and separated oscillating domains for increasingly solid-like systems. Moreover, we illustrate how tuning the rheological properties in such active suspensions can be exploited for active collective transport on the microscopic scale. This tuning allows for intermediate storage times, maintaining the spatial arrangement within a set of cargo units.
@@
In short, in the theoretical framework, the behavior is governed by the interplay of three continuum fields [44]. The coarse-grained polar order parameter field $\mathbf{P}$ characterizes the local alignment of active units. It couples to the overall velocity field $\mathbf{v}$ and to the memory field of elastic displacements $\mathbf{u}$. The latter captures the currently memorized deformation history of the film.
@@
where $\mathbf{I}$ is the unit matrix. In addition to advection via the flow field $\mathbf{v}$, Eq. (1) includes reorientation via the vorticity tensor $\boldsymbol{\Omega}=[(\nabla\mathbf{v})^{\top}-(\nabla\mathbf{v})]$, where denotes the transpose. The first two terms on the right-hand side arise due to orientational and translational diffusion, while the diffusion coefficients were used to rescale time and space. The term proportional to the alignment strength $\gamma_{\mathrm{a}}$ captures the tendency of active units to align with the overall flow field $\mathbf{v}$ near a supporting substrate [45, 46, 47, 48]. Similar terms are included in various discretized models of active solids [49, 50, 31, 51]. Nonlinearity limits the magnitude of the polar order parameter to $|\mathbf{P}|\leq 1$.
@@
We here assume an incompressible material, $\nabla\cdot\mathbf{v}=0$, enforced by the pressure field $p$. Besides hydrodynamic viscous stress, the equation includes elastic stresses due to previous strain deformations of the material [52, 53, 54, 44]. Viscosity $\eta$ and shear modulus $\mu$ together with the elastic memory field $\mathbf{u}$ quantify these viscous and elastic stresses. Underlying friction with the substrate is set by $\nu_{\mathrm{v}}$ and slows down movement. Similarly, elastic restoring forces quantified by $\nu_{\mathrm{d}}$ result from interactions with the substrate and pull material elements back to their memorized positions. Active forcing drives the dynamics of the film and is set by $\nu_{\mathrm{p}}$.
Third, the dynamics of $\mathbf{u}$ is governed by
@@
These dynamics result systematically for elastic strain in established generalized hydrodynamic theories that include elasticity [55, 52, 53, 54, 44]. From the hydrodynamic Eulerian perspective, $\mathbf{u}$ contains the memory of the locations from where the material elements have been displaced to their current positions. We therefore rather call it memory field instead of displacement field [56]. For viscoelastic materials, this memory may fade away over time [52, 53, 54, 44]. On the right-hand side of Eq. (3) overall motion $\mathbf{v}$ generates displacements and thus associated memory. Conversely, the memory of displacement $\mathbf{u}$ decays with a relaxation time $\tau_{\mathrm{d}}$ for incomplete elasticity. This relaxation process implies a net flow. Full elastic reversibility for solids is found in the limit $\tau_{\mathrm{d}}\to\infty$. Finite values $\tau_{\mathrm{d}}$ describe a broad range of viscoelastic fluids, while the limit $\tau_{\mathrm{d}}\to 0$ refers to purely viscous fluids. Thus, rheology can be tuned continuously from fluid- to solid-like behavior via $\tau_{\mathrm{d}}$. Equation (3) includes convective transport $\mathbf{v}\cdot\nabla\mathbf{u}$ by overall flow. In combination with Eq. (2), it represents a first step beyond linear elasticity [56]. This description can be derived from a generic continuum theory [52] based purely on conservation laws and symmetry arguments upon linearization in $\nabla\mathbf{u}$ [44]. The contribution $-\nabla q$ arises from incompressibility $\nabla\cdot\mathbf{u}=0$ under linearized strains [44].
Figure 1: Dynamical state diagram as a function of the displacement relaxation time $\tau_{\mathrm{d}}$ and strength of active forcing $\nu_{\mathrm{p}}$. From left to right, the influence of elasticity increases. Colors indicate regions of different stable dynamical states. Insets in the corners illustrate idealized tracer trajectories. Blue and red dots refer to the dynamical states explored in more detail in Figs. 2 and 3, respectively. Remaining parameter values are $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, and $\nu_{\mathrm{d}}=10$.
Previous derivation of the theory suggested three idealized, analytical, spatially uniform mathematical solutions [44]. We here challenge their stability, amended by numerical simulations. For technical details see the Supplemental Material [57], which includes Refs. 58, 59, 60. The central goal is to construct the dynamical state diagram for active films of viscoelastic rheology.
@@
Raising the influence of elasticity by increasing $\tau_{\mathrm{d}}$, we find that both states extend to the right in Fig. 1. Yet, elasticity seems to hinder motion. Illustratively, the elastic restoring forces pin the active agents. Therefore, the phase boundary between polar moving (green) and isotropic resting (yellow) areas turns towards higher activity with increasing elasticity. The phase boundary results from linear stability analysis and is confirmed numerically [57].
The picture changes notably when we increase elasticity at elevated levels of activity from the polar state. Then novel, qualitatively different behavior emerges. It is marked by the blue region in Fig. 1. Numerical evaluations indicate a stripe-like dynamic state, see Fig. 2(a). The inset depicts associated wiggling, snake-like motion.
@@
Indeed, the state of globally uniform polar orientational order becomes linearly unstable in this region [57]. For the associated linear stability analysis [57], we summarize all dynamic quantities in a vector $\mathbf{V}$. We consider small perturbations $\delta\mathbf{V}$ in terms of Fourier modes of wavevector $\mathbf{k}$, $\delta\mathbf{V}=\delta\hat{\mathbf{V}}e^{\lambda t}e^{i\mathbf{k}\cdot\mathbf{x}}$. The growth rate $\mathrm{Re}\{\lambda(k=|\mathbf{k}|)\}$ is plotted in Fig. 2(b), with $\mathbf{k}$ oriented parallel to $\mathbf{P}$. Importantly, the fastest-growing mode is of finite wavenumber $k_{\mathrm{max}}>0$. Thus, the instability features a certain length scale $\Lambda=2\pi/k_{\mathrm{max}}$, at least close to its onset. This length is reflected by spatial modulations along the migration direction, see Fig. 2(a). There, the local velocity field is indicated by the dark arrows. Its directions deviate periodically in space from the average migration direction given by $\langle\mathbf{v}\rangle$. The periodic modulation is further visualized by the vorticity field $\omega=(\nabla\times\mathbf{v})_{z}$, see the colored stripe pattern in Fig. 2(a).
@@
Moreover, from $\mathrm{Im}\{\lambda(k)\}$, we calculate the phase velocity of the pattern, $v_{\mathrm{ph}}=-\mathrm{Im}(\lambda)/k$. It characterizes the traveling speed of the individual waves. Interestingly, we always find $v_{\mathrm{ph}}<0$. Consequently, the emerging structures travel in the direction opposite to global motion $\langle\mathbf{v}\rangle$, see Fig. 2(a), see also Supplemental Movie 1.
Banded structures are observed in various systems of active matter. Examples include traveling density bands in systems of self-propelled particles [61], bands in suspensions of microswimmers subject to external fields [62], as well as active nematic systems [63]. In contrast to these earlier observations, the traveling band-like structures we find in our study at the boundary of the ordered polar state are induced by elasticity, only observable in the viscoelastic regime.
@@
Figure 3: Heterogeneous state of imperfect local circling motion. (a) At each spot, the vectors of polar orientational order $\mathbf{P}$ and collective flow $\mathbf{v}$ (dark arrows) rotate over time. Domains are identified by the field of phase shift $\Delta\phi_{Pv}$ (color bar) between $\mathbf{P}$ and $\mathbf{v}$. Positive and negative values are connected to clockwise and counterclockwise rotation, respectively. A sample trajectory is shown in the top right corner. The size of the snapshot is $100\times 100$. (b) Averaged spatial correlation function $C_{\Delta\phi_{Pv}}(\Delta x)$ of the phase shift $\Delta\phi_{Pv}$ between any two points of evaluation with distance $\Delta x$ from each other. The associated characteristic length scale $\ell$ of the dynamic patterns is identified with the first minimum of $C_{\Delta\phi_{Pv}}(\Delta x)$. (c) Typical domain size $\ell$ as a function of active force strength $\nu_{\mathrm{p}}$ for different values of the relaxation time $\tau_{\mathrm{d}}$. (d) Enlarged view of the marked area in (a) showing $+1$ and $-1$ defects in the velocity field as red and blue dots. (e) Time-averaged number of defect pairs $\langle N_{v}\rangle$ in the velocity field of a system of size $200\times 200$ as a function of $\nu_{\mathrm{p}}$. Remaining parameter values are $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, $\nu_{\mathrm{d}}=10$, and $\nu_{\mathrm{p}}=20$.
Further increasing the influence of elasticity via $\tau_{\mathrm{d}}$ to the right in Fig. 1, we cross the red line. It marks the existence of a basic analytical solution, corresponding to rotating spatially uniform polar order $\mathbf{P}$, flow $\mathbf{v}$, and elastic memory $\mathbf{u}$ [57]. Yet, we do not observe this state in our spatially extended numerical systems.
Here, numerically, we identify a hysteretic regime until we reach the dotted line in Fig. 1 [57]. In this area, the emergent state depends sensitively on the initial conditions. On the one hand, upon initialization by a weakly perturbed state of global polar orientational order, still the stripe patterns develop. On the other hand, upon initialization by a weak perturbation of the above-mentioned rotational analytical solution [57], rotations of the three fields are maintained. Further to the right of the dotted line in Fig. 1, the stripes are not found as a final dynamic state any more. Although transients can be observed for different initializations, at the end, we find a state of locally rotating fields $\mathbf{P}$, $\mathbf{v}$, and $\mathbf{u}$, see Fig. 3(a). In fact, this state is not spatially uniform. The system splits up into different domains of rotation. Thus, it becomes spatially heterogeneous.
Physically, the elastic restoring forces for elevated $\tau_{\mathrm{d}}$ become too strong to allow for persistent motion. The globally ordered polar or stripe-like dynamic states are associated with continuous directed transport. Thus, they are suppressed by higher levels of elasticity. Instead, local rotations of the polar order $\mathbf{P}$, flow field $\mathbf{v}$, and memory field $\mathbf{u}$ combine active transport with local elastic anchoring. Imperfect circular trajectories arise, see the top right insets in Figs. 1 and 3(a).
@@
Here, $\Delta x$ is the distance between any two points in space and $\hat{\mathbf{n}}$ is the direction of their separation. $\langle\dots\rangle_{\hat{\mathbf{n}},\mathbf{x},t}$ denotes the average over orientations $\hat{\mathbf{n}}$, space positions $\mathbf{x}$, and time $t$. For the dynamically evolving state depicted in Fig. 3(a), we plot $C_{\Delta\phi_{Pv}}(\Delta x)$ in Fig. 3(b). The distinct length scale $\ell$ quantifying the typical domain size is identified with the location of the first minimum of $C_{\Delta\phi_{pv}}(\Delta x)$. Indeed, we find that the length scale $\ell$ of the typical size of the correlated domains is of the same order as the thickness of the stripes $\Lambda/2$ referred to in Fig. 2(a). Figure. 3(c) shows $\ell$ as a function of the strength of active forcing $\nu_{\mathrm{p}}$ for different values of the relaxation time $\tau_{\mathrm{d}}$ within the red region of Fig. 1. Mainly, we find that $\ell$ obtains a minimum, but then increases continuously with increasing activity $\nu_{\mathrm{p}}$. Thus, the structures become increasingly correlated. Approaching the transition to the isotropic state, see below, $\ell$ diverges and the structures are highly correlated in space. Overall, we find that the relaxation time of elasticity in terms of $\tau_{\mathrm{d}}$ has only little influence on $\ell$ for the considered parameters.
As for various other systems of active matter [64, 9, 22, 65], defects in the velocity and polar order parameter fields are continuously pairwise created and annihilated. Their mobility is visualized in Supplemental Movie 3. In contrast to active nematics [64, 9], we observe defects of integer value in our polar system. $+1$ defects correspond to vortices, whereas $-1$ defects locate (saddle) points in between, see Fig. 3(d). We expect the number of defects to behave oppositely to the domain size $\ell$ in Fig. 3(c). Indeed, Fig. 3(e) confirms this relation for the time-averaged number of defect pairs $\langle N_{v}\rangle$ in the velocity field as a function of activity $\nu_{\mathrm{p}}$ for different values of $\tau_{\mathrm{d}}$. The time-averaged number of defect pairs for the field $\mathbf{P}$ yields qualitatively the same result [57].
@@
As an immediate perspective, changing the rheology of active films allows to control collective transport. For example, light and external electric fields enable remote dynamic tuning of the rheology of photorheological [66, 67, 68] or electrorheological [69] materials. Exploiting such effects facilitates externally controlled transport of cargo objects by active carrier media. Persistent polar orientational order implies directed transport, while switching to the nonpolar states facilitates intermediate storage times. Elevated activity implies increased speed of transport. It requires intermediate switching to the state of imperfect local circling during rest intervals. We numerically demonstrate this scenario in Fig. 4 by periodically changing the relaxation time $\tau_{\mathrm{d}}$, see Fig. 4(a). Resulting trajectories in Fig. 4(b) show the transport of an array of cargo objects, including intermediate storage periods, while maintaining the internal positional order within the array. Suitable preparation of the substrate can guide the transport. The approach opens new routes in microfluidics powered by the activity of microswimmers such as bacteria [70, 71, 72, 73]. Instead varying the relaxation time in space suggests a potential mechanism for persistent particle trapping.
In summary, we uncover a rich world of dynamic states that result from the interplay between viscoelasticity and activity. We construct the overall dynamical state diagram and observe the emergence of novel dynamic states. First, the stripe-like patterns are neither found for purely viscous active fluids nor for purely elastic active solids. Their origin is thus different from earlier observed banded structures [61, 63, 62]. Second, the heterogeneous states of imperfect circling result from elastic forces within the system, in contrast to heterogeneous states observed in purely viscous suspensions of microswimmers [22, 65].
Besides suggesting engineered active systems to control directed transport of cargo on the microscale, our description relates to the complex dynamical behavior in living biological matter. Biofilms and similar collections of cells on substrates show complex flow patterns [74, 75], which are reminiscent of the structures observed in Fig. 3(a). Bacterial biofilms [2, 3, 4], in particular, may change their rheological characteristics dynamically and continuously from purely viscous via viscoelastic to elastic as a function of time and position within the film.
@@
Figure 4: Rheologically controlled transport of an array of cargo objects. (a) We periodically vary the relaxation time $\tau_{\mathrm{d}}$ of elastic displacements between $0.25$ and $1$. Thus, we tune the degree of elasticity. (b) Resulting trajectories for an array of four transported cargo objects. Tuning rheology allows in a controlled way to switch between persistent collective transport and intermediate storage times, roughly maintaining the spatial order within the transported cargo array. Blue and red parts in (a) and (b) correspond to each other and to the blue and red regions in the state diagram in Fig. 1. The remaining parameters are $\gamma_{\mathrm{a}}=1$, $\eta=1$, $\mu=1$, $\nu_{\mathrm{v}}=1$, $\nu_{\mathrm{d}}=10$, and $\nu_{\mathrm{p}}=20$. The initial edges in the cargo array are of length $1$.